OpenStudy (anonymous):

Calculus II: Exponential Change Problem? Can you please help me with this problem: The Charcoal From tree killed in the volcanic eruption that formed Crater Lake in Oregon contained 44.5% of the carbon-14 found in living matter. About how old in Crater Lake?

OpenStudy (bahrom7893):

The fraction of original amount present in a sample is µ = y/y₀ = e^(- k·t ) <=> t = - (1/k)·ln(µ) = - 8266y · ln( 0.445 ) = 6693y

OpenStudy (anonymous):

where did that( - 8266y) come from :/

OpenStudy (bahrom7893):

okay so: t(1/2) = 5730 y <= half life of C14

OpenStudy (bahrom7893):

1/k = t1/2 / ln(2) = 8266y

OpenStudy (bahrom7893):

1/k = (t[1/2]) / ln(2) = 8266y

OpenStudy (anonymous):

oh ok i see... but in the book..it says the answer is 6658 years..does it matter?

OpenStudy (bahrom7893):

no, half life of C14 is not very precise... we'll always be off by several years, that's not really relevant

OpenStudy (bahrom7893):

depends on how u round off, etc.. we are off by how many around 35 years. I know it seems a lot, but i think its fine. Fan me if I helped. Thanks

OpenStudy (bahrom7893):

http://faculty.eicc.edu/bwood/ma155supplemental/supplemental5.htm btw just found ur exact problem there. they are getting ur book's answer, u might wanna take a look at example 4

OpenStudy (anonymous):

Thank you so much! how do i fan you?

OpenStudy (bahrom7893):

Click on Become a fan next to my name lol